# Discerning What Matters: A Multi-Dimensional Assessment of Moral Competence in LLMs

### v1

@@ 1.1 Background & motivation
I. Over-reliance on pre-packaged cases: The overwhelming majority of ethical evaluations rely on static datasets with neatly prepackaged moral scenarios (Hendrycks et al.,, 2023; Jin et al.,, 2022; Lourie et al.,, 2021). These scenarios typically present clear moral dilemmas with explicitly highlighted morally relevant features. While useful for standardization, this approach fails to assess a model’s ability to identify morally relevant features from noisy, unfiltered information — a capacity we call moral sensitivity. But, as we show below, pre-packaging a scenario means doing much of the work for the model, and then only evaluating the model’s performance on what remains. In real-world ethical decision-making, both human moral advisors and autonomous agents must first identify which features of a situation carry moral significance before rendering judgment.
II. Focus on verdicts over reasoning: Current evaluations of LLM ethical capabilities fundamentally conflate two distinct tasks: predicting human moral judgments versus engaging in sound moral reasoning. This distinction is crucial. Consider mathematical reasoning as an analogy: a system that perfectly predicts how humans answer mathematical questions would not demonstrate mathematical competence. Such a system might excel at reproducing common human errors—like believing division always makes numbers smaller or that multiplication always increases quantities—without understanding mathematical principles. Similarly, systems like DELPHI (Jiang et al.,, 2022) effectively model common ethical intuitions but fail to demonstrate the deeper competence required for ethical reasoning. Albrecht et al.,, 2022 empirically demonstrated this limitation, showing that language models can achieve “superhuman” performance on ethics datasets while making fundamental moral errors humans rarely make. More troublingly, these researchers could flip models’ ethical verdicts through simple rewording that preserved moral content—revealing reliance on linguistic patterns rather than moral understanding. This gap between predicting popular verdicts and demonstrating genuine moral reasoning ability represents a critical blind spot in evaluating whether LLMs could serve as reliable moral cores for autonomous systems increasingly impacting human lives.
@@ 2.1 Materials
LLM vignette responses. We prompted each of the six LLMs to respond to all 12 vignettes via API calls and responses were collected in the same JSON format from all models. We used default values of each model for temperature (randomness of model output), maximum token and other settings. We presented each vignette as a new API call with the same system prompt, and a user prompt containing the vignette and the same 5 questions in the human survey. [^1] System and user prompt details are included in Appendix 5 and the Python code (blinded for peer review) can be found on https://anonymous.4open.science/r/Measuring-Moral-Skill-in-LLMs-BE23
Initially, we generated five responses per model, but since each LLM’s responses were largely self-similar, we used only the first response from each model for each vignette. To quantify the semantic similarity (cosine similarity), we used DistilRoBERTa transformer-based language model as a cross encoder (Sanh et al.,, 2020); average scores per model ranged from 0.58 to 0.82 (1 indicates very similar semantic properties).
@@ 2.2 Human Assessors
We recruited 590 participants333599 were initially recruited, however seven participants were excluded because they did not agree to have their data anonymised and two were excluded because they selected only the first or only the second option throughout the study, resulting in a final sample of 590 participants. No participants were excluded for failing an attention check, or for taking less than 30 seconds to complete the survey. via Prolific to provide comparative judgements on the quality of pairs of vignette responses. Half of the participants were female, 48% male, 2% identified as non-binary, and 1% did not disclose gender. All age categories from 18 to 84 were represented with 55% of participants falling in the 25-44 age group. The majority (69%) of participants reside in the United States of America and Canada. 47% declared no religion and 38% were Christians, the remaining 15% of participants follow other religions. Almost all participants have at least high school or higher education and 42% hold a bachelors or undergraduate degree. Full demographic details are reported in Appendix 2.
@@ 2.5 Results
Figure 1: Bradley–Terry model estimates of LLM performance identifying morally salient features in published vignettes (Question 1). The dotted line indicates the performance of the general public (human reference group). Error bars show 95% confidence intervals.
[…]
Figure 2: Bradley–Terry model estimates of LLM performance on identifying additional relevant information for novel vignettes (Question 5)
[…]
@@ 3.1 Materials
Vignettes. We created 12 entirely novel moral scenarios drawn from an unpublished database of naturalistic vignettes. Our methodical approach to vignette creation involved selecting 30 third-person narratives from a pool of 351, replacing generic names with ethnically diverse alternatives. From these modified scenarios, an independent expert in moral philosophy selected the best 12 cases, establishing face validity of our survey (see Appendix 3). These were specifically designed to eliminate the risk of training data contamination for the LLMs. When developing these vignettes, we employed three specific strategies:
@@
    Deliberately including morally irrelevant information to test discernment abilities
  3. 3.
@@
LLM vignette responses. The procedure for collecting LLM response was the same as in our first experiment.
@@ 3.2 Human Assessors
We recruited 586 participants444600 were initially recruited, however three were excluded for not agreeing to having their data anonymised, and eleven were excluded for selecting only the first or only the second option throughout the study, resulting in a final sample of 586 participants. As in Study 1, no participants completed the survey in less than 30 seconds or failed an attention check qeuestion. to serve as human assessors via Prolific. This experiment attracted more female participants (60%) than the first experiment. Age profile is similar with 55% of participants falling in the 25-44 age group. Unlike the first experiment, the majority (73%) of participants reside in the United Kingdom. Similar to the first experiment, 51% declared no religion, 39% were Christians, with the remaining 10% of participants following other religions. Education profile was very similar to the first experiment with all participants having at least high school or higher education and 42% holding a bachelor’s / undergraduate degree. Full demographic details are reported in Appendix 2.
@@ 3.4 Results
[…]
@@ 3.5.1 Limitations
Finally, our experiments were not designed to investigate correlations between moral reasoning subtasks. Determining whether these skills operate independently or together would provide valuable insights into the structure of moral competence. Our focus on relative rankings rather than absolute metrics prevented such correlation analysis, and future experiments could relax this limitation.
@@ Acknowledgments
Gender Number of participants Percentage Number of participants Percentage
@@
Age Number of participants Percentage Number of participants Percentage
@@
Country of Residence Number of participants Percentage Number of participants Percentage
@@
Religion Number of participants Percentage Number of participants Percentage
@@
Education Number of participants Percentage Number of participants Percentage
@@
The dotted line indicates the performance of the general public (human reference group). Error bars show 95% confidence intervals.
Figure 9: Experiment 1 Bradley–Terry model estimates of systems’ performance on identifying morally salient features in published vignettes (Question 1)
Figure 10: Experiment 2 Bradley–Terry model estimates of systems’ performance on identifying morally salient features in published vignettes (Question 1)
Figure 11: Experiment 1 Bradley–Terry model estimates of systems’ performance on weighting morally salient features in novel vignettes (Question 2)
Figure 12: Experiment 2 Bradley–Terry model estimates of systems’ performance on weighting morally salient features in novel vignettes (Question 2)
Figure 13: Experiment 1 Bradley–Terry model estimates of systems’ performance on associating reasons with morally salient features in novel vignettes (Question 3)
Figure 14: Experiment 2 Bradley–Terry model estimates of systems’ performance on associating reasons with morally salient features in novel vignettes (Question 3)
Figure 15: Experiment 1 Bradley–Terry model estimates of systems’ performance on giving clear moral conclusions for novel vignettes (Question 4)
Figure 16: Experiment 2 Bradley–Terry model estimates of systems’ performance on giving clear moral conclusions for novel vignettes (Question 4)
Figure 17: Experiment 1 Bradley–Terry model estimates of systems’ performance on identifying additional relevant information for novel vignettes (Question 5)
Figure 18: Experiment 2 Bradley–Terry model estimates of systems’ performance on identifying additional relevant information for novel vignettes (Question 5)

### v3

@@ 1.1 Background & motivation
I. Over-reliance on pre-packaged cases: The overwhelming majority of ethical evaluations rely on static datasets with neatly prepackaged moral scenarios (Hendrycks et al.,, 2023; Jin et al.,, 2022; Lourie et al.,, 2021), often derived from classic ethical thought experiments (Jin et al.,, 2024; bin Ahmad and Takemoto,, 2024) or from psychological instruments (Nunes et al.,, 2023; Tanmay et al.,, 2023). These scenarios typically present clear moral dilemmas with explicitly highlighted morally relevant features. While useful for standardization, this approach fails to assess a model’s ability to identify morally relevant features from noisy, unfiltered information — a capacity we call moral sensitivity. But, as we show below, pre-packaging a scenario means doing much of the work for the model, and then only evaluating the model’s performance on what remains. In real-world ethical decision-making, both human moral advisors and autonomous agents must first identify which features of a situation carry moral significance before rendering judgment.
II. Focus on verdicts over reasoning: Current evaluations of LLM ethical capabilities fundamentally conflate two distinct tasks: predicting human moral judgments versus engaging in sound moral reasoning. For instance, many LLM ethical evaluations focus on judgment prediction with human situational ethics data serving as the gold standard (Bignotti and Camassa,, 2024; Bulla et al.,, 2025) such as the ‘Am I The Asshole’ subreddit (Russo et al.,, 2025; Sachdeva and van Nuenen,, 2025). But the prediction vs. reasoning distinction is crucial. Consider mathematical reasoning as an analogy: a system that perfectly predicts how humans answer mathematical questions would not demonstrate mathematical competence. Such a system might excel at reproducing common human errors—like believing division always makes numbers smaller or that multiplication always increases quantities—without understanding mathematical principles. Similarly, systems like DELPHI (Jiang et al.,, 2022) effectively model common ethical intuitions but fail to demonstrate the deeper competence required for ethical reasoning. Albrecht et al.,, 2022 empirically demonstrated this limitation, showing that language models can achieve “superhuman” performance on ethics datasets while making fundamental moral errors humans rarely make. More troublingly, these researchers could flip models’ ethical verdicts through simple rewording that preserved moral content—revealing reliance on linguistic patterns rather than moral understanding. This gap between predicting popular verdicts and demonstrating genuine moral reasoning ability represents a critical blind spot in evaluating whether LLMs could serve as reliable moral cores for autonomous systems increasingly impacting human lives.
@@ 2.1 Materials
LLM vignette responses. We prompted each of the six LLMs to respond to all 12 vignettes via API calls and responses were collected in the same JSON format from all models. We used default values of each model for temperature (randomness of model output), maximum token and other settings. We presented each vignette as a new API call with the same system prompt, and a user prompt containing the vignette and the same 5 questions in the human survey [^1] System and user prompt details are included in Appendix 5 and the Python code (blinded for peer review) can be found on https://anonymous.4open.science/r/Measuring-Moral-Skill-in-LLMs-BE23. The second question asked to rank the features by assigning each feature a weight and that all weights must add up to 100 for each vignette. We observed that all answers to this question were as instructed.
Initially, we generated five responses per model, but since each LLM’s responses were largely self-similar, we used only the first response from each model for each vignette [^2] All responses are available on https://anonymous.4open.science/r/Measuring-Moral-Skill-in-LLMs-BE23/llm_answers_experiment1.md. To quantify the semantic similarity (cosine similarity), we used DistilRoBERTa transformer-based language model as a cross encoder (Sanh et al.,, 2020); average scores per model ranged from 0.58 to 0.82 (1 indicates very similar semantic properties).
@@ 2.2 Human Assessors
We recruited 590 participants444599 were initially recruited, however seven participants were excluded because they did not agree to have their data anonymised and two were excluded because they selected only the first or only the second option throughout the study, resulting in a final sample of 590 participants. No participants were excluded for failing an attention check, or for taking less than 30 seconds to complete the survey. via Prolific to provide comparative judgements on the quality of pairs of vignette responses. Half of the participants were female, 48% male, 2% identified as non-binary, and 1% did not disclose gender. All age categories from 18 to 84 were represented with 55% of participants falling in the 25-44 age group. The majority (69%) of participants reside in the United States of America and Canada. 47% declared no religion and 38% were Christians, the remaining 15% of participants follow other religions. Almost all participants have at least high school or higher education and 42% hold a bachelors or undergraduate degree. Full demographic details are reported in Appendix 2.
@@ 2.5 Results
Figure 1: Bradley–Terry model estimates of LLM performance identifying morally salient features in published vignettes (Question 1 on the left and Question 5 on the right). The dotted line at 0 indicates the performance of the general public (human reference group). Error bars show 95% confidence intervals.
Identification of morally salient features: For the first question, four LLMs significantly outperformed human responses $(\alpha=0.05)$: GPT-4o $(\beta=0.33,p<.001)$, GPT-o1 $(\beta=0.24,p<.001)$, and DeepSeek Chat $(\beta=0.17,p<.01)$. Claude 3.7 Sonnet also performed significantly better than humans $(\beta=0.14,p=.03)$. The remaining LLMs—Llama 3.3 and Gemini 1.5 Flash—did not differ significantly from humans in their ability to identify morally salient features.
Weighting moral features: When assessing the ability to appropriately weight the relative importance of identified moral features (Question 2), three LLMs again demonstrated superior performance compared to humans: GPT-4o $(\beta=0.28,p<.001)$, GPT-o1 $(\beta=0.25,p<.001)$, and Claude 3.7 Sonnet $(\beta=0.18,p<.01)$. DeepSeek Chat, Gemini 1.5 Flash, and Llama 3.3 showed no significant difference from human responses.
Associating reasons with features: For Question 3, only GPT-4o significantly outperformed humans $(\beta=0.18,p<.01)$. The remaining five LLMs showed no significant difference from human performance in their ability to associate reasons with identified moral features.
Forming clear moral conclusions: When evaluating the ability to combine reasons into clear moral conclusions (Question 4), only Llama 3.3 significantly outperformed humans $(\beta=0.20,p<.01)$. The other five LLMs did not differ significantly from humans.
Identifying additional relevant information: For the final question, four LLMs significantly outperformed humans: DeepSeek Chat $(\beta=0.22,p<.001)$, Llama 3.3 $(\beta=0.20,p<.01)$, GPT-4o $(\beta=0.17,p<.01)$, and Gemini 1.5 Flash $(\beta=0.16,p<.05)$. Claude 3.7 Sonnet and GPT-o1 showed no significant difference from humans.
@@
To measure the consistency of the human judgements of the responses, we calculated per-item agreement. This is the proportion of raters who, for each specific vignette × question × system‑pair, chose the majority option. The mean item agreement for the first study was 0.724 (SD = 0.144), indicating that, on average, nearly three‑quarters of raters concurred on the “winner.” A Monte Carlo simulation under random selection confirmed that this exceeds chance $p<.001$, suggesting good consistency in people’s judgements of which responses were best.
@@ 3.1 Materials
Vignettes. We created 12 entirely novel moral scenarios drawn from an unpublished database of naturalistic vignettes. When developing these vignettes, we employed three specific strategies:
@@
Our methodical approach to vignette creation involved selecting 30 third-person narratives from a pool of 351, replacing generic names with ethnically diverse alternatives. We have then asked GPT-4o to modify them according to strategies 1 and 2  [^5] The prompts used to modify the vignettes can be found in https://anonymous.4open.science/r/Measuring-Moral-Skill-in-LLMs-BE23/modify_vignettes.py. From these modified scenarios, an independent expert in moral philosophy selected the best 12 cases, establishing face validity of our survey (see Appendix 3). These were specifically designed to eliminate the risk of training data contamination for the LLMs.
@@
LLM vignette responses. The procedure for collecting LLM response was the same as in our first experiment [^6] All responses are available on https://anonymous.4open.science/r/Measuring-Moral-Skill-in-LLMs-BE23/llm_answers_experiment2.md. We again observed that all answers to the second question of weighing features added up to 100.
@@ 3.2 Human Assessors
We recruited 586 participants777600 were initially recruited, however three were excluded for not agreeing to having their data anonymised, and eleven were excluded for selecting only the first or only the second option throughout the study, resulting in a final sample of 586 participants. As in Study 1, no participants completed the survey in less than 30 seconds or failed an attention check qeuestion. to serve as human assessors via Prolific. This experiment attracted more female participants (60%) than the first experiment. Age profile is similar with 55% of participants falling in the 25-44 age group. Unlike the first experiment, the majority (73%) of participants reside in the United Kingdom. Similar to the first experiment, 51% declared no religion, 39% were Christians, with the remaining 10% of participants following other religions. Education profile was very similar to the first experiment with all participants having at least high school or higher education and 42% holding a bachelor’s / undergraduate degree. Full demographic details are reported in Appendix 2.
@@ 3.4 Results
Figure 2: Bradley–Terry model estimates of systems’ performance on identifying morally salient features in novel vignettes (Question 1 on the left and Question 2 on the right). The dotted line at 0 indicates the performance of the general public (human reference group). Error bars show 95% confidence intervals.
Identification of morally salient features: For the first question, three LLMs performed significantly worse than non-philosopher responses: Claude 3.7 Sonnet ($\beta=-0.35$, $p<.001$), Gemini 1.5 Flash ($\beta=-0.21$, $p<.01$), and DeepSeek Chat ($\beta=-0.14,p<.05$). The remaining systems—Llama 3.3, GPT-4o, GPT-o1, and Philosopher responses—did not differ significantly from human responses in identifying morally salient features.
Weighting moral features: When assessing the ability to appropriately weight the relative importance of identified moral features (Question 2), two LLMs demonstrated significantly worse performance compared to humans: Claude 3.7 Sonnet $(\beta=-0.34,p<.001)$ and Gemini 1.5 Flash $(\beta=-0.17,p<.05)$. DeepSeek Chat showed marginally worse performance $(\beta=-0.12,p<.10)$. The remaining systems—GPT-4o, Llama 3.3, GPT-o1, and Philosopher responses—showed no significant difference from human responses.
@@
For this second study, the per-item agreement was 0.734 (SD = 0.148), indicating once again that, on average, nearly three‑quarters of raters concurred on the “winner” of a pairwise comparison. Monte Carlo simulation under random selection again confirms that this exceeds chance $p<.001$.
@@ 4 Limitations
Third, our experiments were not designed to investigate correlations between moral reasoning subtasks. Determining whether these skills operate independently or together would provide valuable insights into the structure of moral competence. Our focus on relative rankings rather than absolute metrics prevented such correlation analysis, and future experiments could relax this limitation.
Finally, our vignette set comprises 12 cases that do not span the full moral domain; they centre on interpersonal dilemmas. Though this represents an important limitation to our study, we made this restriction intentionally: first, to enhance ecological validity for a key motivating application—Artificial Moral Advisors that assist with everyday, interpersonal decisions—so evaluating that surface first is appropriate; second, to avoid muddying the analysis with legal, institutional, or political-legitimacy questions that risk conflating moral assessment with other flavours of normativity; and third, to maximise comparability with prior work, which predominantly uses short, person-centred scenarios. This continuity let us cleanly demonstrate the “reversal” we observe when superficial cues are removed.
@@ Acknowledgments
Gender
Number of
participants
Percentage
Number of
participants
Percentage
@@
Age
Number of
participants
Percentage
Number of
participants
Percentage
@@
Country of Residence
Number of
participants
Percentage
Number of
participants
Percentage
@@
Religion
Number of
participants
Percentage
Number of
participants
Percentage
@@
Education
Number of
participants
Percentage
Number of
participants
Percentage
@@
The dotted line at 0 indicates the performance of the general public (human reference group). Error bars show 95% confidence intervals. Experiment 1 results are shown on the left and Experiment 2 results on the right.
Figure 9: Bradley–Terry model estimates of systems’ performance on associating reasons with morally salient features in novel vignettes (Question 3)
Figure 10: Bradley–Terry model estimates of systems’ performance on giving clear moral conclusions for novel vignettes (Question 4)
Figure 11: Bradley–Terry model estimates of systems’ performance on identifying additional relevant information for novel vignettes (Question 5)